Calculate the WACC for the firm based on this information:

Outstanding shares = 5000 , Share price = $54 , Total debt = 30,000 , Risk-free rate = 2% , Return on the market = 7% , Yield to Maturity = 5.4% , Tax rate = 0.25 , Beta = 0.55
Answer #1

In order to calculate WACC we will use the formula : (E/T.V.)*k(e) + (D/T.V.)*(1-T)*(k(d))

where, E = Market value of Equity D = Market value of debt

T.V. = Total market value of Equity and debt k(d) = Cost of debt

k(e) = Cost of equity T = Corporate tax rate

Here all the information is given in the question except cost of equity k(e), so we will calculate cost of equity using CAPM:

k(e) = R(f) + ( R(m) - R(f) )*Beta

where, R(f) = Risk-free rate R(m) = Return on the market

So, k(e) = 0.02 + (0.07 - 0.02 ) * 0.55

= 0.0475 or 4.75%

Now, WACC :

WACC = (E/T.V.)*k(e) + (D/T.V.)*(1-T)*(k(d))

E = 5000*54 = 270000 D = 30000

T.V. = 270000 + 30000 = 300000 k(e) = 0.0475 k(d) = Yield to maturity = 5.45% = 0.0545

WACC = ( 270000 / 300000)*(0.0475) + (30000/300000)*(1-0.25)*(0.0545)

= 0.04275 + 0.0040875

= 0.0468 or 4.68%

Therefore, WACC = 4.68%